Peanuts cost $0.80 a pound, cashews $1.50 a pound, anf alnuts $1.50 a pound. A store clerk needs to make 6 pound of mixed nuts using these three types so the cost is $6.90. If she wants to use 3 times as much peanuts as cashews, find how much of each type she should use.
Answer by checkley79(3341)

.80P+1.50C+1.50A=6*6.90
P=3C
A=C
.80(3C)+1.50C+1.50C=41.40
2.4C+3C=41.40
5.4C=41.40
C=41.40/5.4
C=7.67 POUNDS OF CASHEWS & ALNUTS.
P=3*7.67=23 POUNDS OF PEANUTS.
PROOF:
.80*23+1.5*7.67+1.5*7.67=6*6.90
18.40+11.5+11.5=41.40
41.40=41.40